What is the formula for measuring the price elasticity of supply? Suppose the price of apples goes up from $20 to $22 a box. In direct response, Goldsboro Farms supplies 1,200 boxes of apples instead of 1,000 boxes. Compute the coefficient of price elasticity (midpoint approach) for Goldsboro’s supply. Is its supply elastic, or is it inelastic?


1
Expert's answer
2021-11-09T13:25:34-0500

Formula for price elasticity of supply:

Price Elasticity of supply= percentage change in quantity supplied/percentage change in price.

"=\frac{Percentage Change in Quantity supplied}{Percentage Change in Price}"


"PES=\frac{\frac{1200-1000}{1100}\times100}{\frac{22-20}{21}\times100}"

"PES=1.91=2."

Since price elasticity of supply is greater than 1, the supply is elastic.